OpenStudy (anonymous):

\(\sf 16x^2 - 24x+7 = 0\)

OpenStudy (abhisar):

Use the quadratic formula, \(\sf x = \huge \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (sepeario):

When using the quadratic formula, remember that you are using a quadratic in the form of ax^2 + bx + c = 0. So in your quadratic that has been given, a=16, b=-24 c=7.

OpenStudy (sepeario):

Can you substitute those values into the given equation above?